A bike is traveling to the left with a speed of 27 (m)/(s). when the rider slams on the brakes. The bike skids for 41.5m with constant acceleration before it comes to a stop. What was the acceleration of the bike as it came to a stop?
1
Expert's answer
2020-07-06T15:13:40-0400

The kinematic law of motion under the constant acceleration is given be the following equation:


"d = v_0t + \dfrac{at^2}{2}"

where "d = 41.5m" is the traveled distance, "v_0 = 27 m\/s" is the initial velocity, "a" is the constant acceleration and "t" is the time of motion.

By definition, the acceleration is given by:


"a = \dfrac{0-v_0 }{t} = -\dfrac{v_0 }{t}"

where 0 stands for the final velocity (bike stops completely after time "t" passes).

Substituting this expression to the law of motion, get:


"d = v_0t - \dfrac{v_0t}{2} = \dfrac{v_0t}{2}"

Expressing "t" and substituting it to the expression for the acceleration, obtain:


"t = \dfrac{2d}{v_0}\\
a = -\dfrac{v_0^2}{2d}"

Calculating the numerical value, get:

"a = -\dfrac{27^2}{2\cdot 41.5}\approx -8.78 m/s^2"

Answer. The acceleration is -8.78 m/s^2.